The percentage of pyridinthat forms pyridinium ion in a 0.10M aqueous solution is
 

Grade:11

1 Answers

Arun
6 years ago
Ka = Kw/Kb = 5.8 x 10^-6 = x^2 / 0.10-x 
x = [C5H5NH+]= 0.00077 M 

% = 0.00077 x 100/ 0.10=0.77
 
 
Regards
Arun (askIITians forum expert)